OpenStudy (anonymous):

Determine if the integral is convergent or divergent. If it is convergent, evaluate. Integral from 0 to 2 1/(x-2)^2

OpenStudy (anonymous):

converges to zero ?

OpenStudy (anonymous):

i would say diverges to -infinity

OpenStudy (anonymous):

it diverges

OpenStudy (dumbcow):

\[\int\limits_{0}^{2}\frac{1}{(x-2)^{2}} = -\frac{1}{x-2} from 0->2\] \[= -\infty - \frac{1}{2}\] thus diverges